If prism has 1.5 refractive index and emission angle is double of incidence angle, what will be the angle of minimum deviation?
1
Expert's answer
2019-07-22T11:37:24-0400


In case of minimum deviation, the angle of deviation D is equal to incidence angle i

"D=i (1)"

Now by Snell’s Law applying at interface of prism:

"\frac { \sin i } {\sin r } = 1.5 (2)"

In our case,


"r=2i (3)"

We put (3) in (2)


"\frac { \sin i } {\sin 2i } = \frac { \sin i } {2 \sin i \cos i } = \frac { 1} {\cos i } = 1.5 (4)"

Using (4) we get:


"\cos i = 0.67 (5)"

"i=48\u00b0 (6)"

We put (6) in (1)


"D=48\u00b0"

The angle of minimum deviation is equal to 48°